OpenStudy (anonymous):

Perform the operation. 8/r^3s^2 (divided by) 6/rs^4 = a) 4s/3r b) 4s^2/3r^2 c) 4s^2/3r^3

OpenStudy (anonymous):

flip the second one, then cancel what you can

OpenStudy (anonymous):

b

OpenStudy (anonymous):

\[\frac{8}{r^3s^2}\times \frac{rs^4}{6}\] is a start

OpenStudy (anonymous):

cancel one \(r\) and 2 \(s\) and reduce \(\frac{8}{6}\) to \(\frac{4}{3}\) and get \[\large \frac{4s^2}{3r^2}\]